find the missing endpoint given one endpoint and midpoint
1)
endpoint- (4,2) midpoint-(6,0)
2)
endpoint- ( -4,1) midpoint-(-1,3)
3)
endpoint-(-3,6) midpoint- (-4,4)
4)
endpoint-(6,-10) midpoint (-1,2)
PLEASE PLEASE PLEASE HELP

Answer by stanbon(48564)

Comment: The coordiinates of a midpoint are always the average of the corresponding coordinates of the endpoints.
1)
endpoint- (4,2) midpoint-(6,0)
Let the endpoint be (x,y)
(4+x)/2=6 and (2+y)/2=0
x = 8 and y = -2
Midpoint is (8,-2)
==========================
2)
endpoint- ( -4,1) midpoint-(-1,3)
Let the midpoint b (x,y)
(x+-4)/2 = -1 and (y+1)/2 = 3
x = 2 and y=5
Midpoint is (2,5)
=======================
Use the same procedure on the following problems.
========================
3)
endpoint-(-3,6) midpoint- (-4,4)
4)
endpoint-(6,-10) midpoint (-1,2)
================
Cheers,
Stan H.